What is the solubility (in g/L to 3 decimal places) of magnesium oxalate (MgC2O4, Ksp= 8.5x10-5) in 0.067 M sodium oxalate?
Answer #1

Na2C2O4 here is Strong electrolyte

It will dissociate completely to give [C2O42-] = 0.067 M

At equilibrium:

MgC2O4 <----> Mg2+ + C2O42-

   s 6.7*10^-2 + s

Ksp = [Mg2+][C2O42-]

8.5*10^-5=(s)*(6.7*10^-2+ s)

Since Ksp is small, s can be ignored as compared to 6.7*10^-2

Above expression thus becomes:

8.5*10^-5=(s)*(6.7*10^-2)

8.5*10^-5= (s) * 6.7*10^-2

s = 1.269*10^-3 M

Molar mass of MgC2O4= 112.33 g/mol

s = 1.269*10^-3 mol/L

To covert it to g/L, multiply it by molar mass

s = 1.269*10^-3 mol/L * 112.33 g/mol

s = 0.143 g/L

Answer: 0.143 g/L

Feel free to comment below if you have any doubts or if this answer do not work